# via Continuized Discrete Diffusion

### v1

@@ Abstract
Vision–Language–Action (VLA) models offer a unified framework for robotic manipulation by integrating visual perception, language understanding, and control generation. Yet existing VLA models still struggle to generalize across diverse tasks, scenes, and camera viewpoints, and often produce coarse or unstable actions. We introduce $\mathcal{E}_{0}$, a continuized discrete diffusion framework that formulates action generation as iterative denoising over quantized action tokens. Compared with continuous diffusion policies, $\mathcal{E}_{0}$ offers two key advantages: (1) discrete action tokens align naturally with the symbolic structure of pretrained VLM/VLA backbones, enabling stronger semantic conditioning; and (2) discrete diffusion matches the true quantized nature of real-world robot control—whose hardware constraints (e.g., encoder resolution, control frequency, actuation latency) inherently discretize continuous signals—and therefore benefits from a Bayes-optimal denoiser that models the correct discrete action distribution, leading to stronger generalization. Compared with discrete autoregressive and mask-based discrete diffusion models, $\mathcal{E}_{0}$ supports a significantly larger and finer-grained action vocabulary and avoids the distributional mismatch introduced by masking-based corruptions—yielding more accurate fine-grained action control. We further introduce a spherical viewpoint perturbation augmentation method to improve robustness to camera shifts without additional data. Experiments on LIBERO, VLABench, and ManiSkill show that $\mathcal{E}_{0}$ achieves state-of-the-art performance across 14 diverse environments, outperforming strong baselines by 10.7% on average. Real-world evaluation on a Franka arm confirms that $\mathcal{E}_{0}$ delivers precise, robust, and transferable manipulation, establishing discrete diffusion as a promising direction for generalizable VLA policy learning.
@@ 1 Introduction
Figure 1: Overview of action modeling paradigms. (a) Discrete modeling: Traditional autoregressive (AR) approaches [brohan2022rt, zitkovich2023rt, kim2024openvla] and recent mask-based discrete diffusion methods [liang2025discrete], which operate over a small discrete action vocabulary. (b) Continuous modeling: Continuous diffusion–based policies [liu2024rdt, xu2025a0] and AR–diffusion hybrids [black2024pi_0, intelligence2025pi05, lin2025onetwovla] that regress continuous actions. (c) Our approach: $\mathcal{E}_{0}$ integrates AR-style conditioning with continuized discrete diffusion, enabling efficient action generation while preserving compatibility with pretrained vision–language backbones and supporting fine-grained action control.
@@
A growing body of work focuses on improving action modeling and generation within VLA frameworks, and recent approaches can be broadly grouped into two families. The first family adopts discrete action parameterizations, including autoregressive (AR) token prediction [brohan2022rt, zitkovich2023rt, kim2024openvla, pertsch2025fast] and mask-based discrete diffusion [nie2025large, wu2025fast, ye2025dream] (see Fig. 1(a)). These models benefit from architectural affinity with pretrained vision–language backbones, but their action vocabularies are typically restricted by language tokenizers, limiting action resolution and hindering fine-grained control. Moreover, mask-based diffusion [nie2025large, wu2025fast, ye2025dream] corrupts tokens by replacing them with masks rather than perturbing them under a principled forward stochastic process. This causes a distributional mismatch that deviates from the true discrete action distribution and breaks forward–reverse consistency, preventing these models from accurately modeling fine-grained actions.
The second family employs continuous diffusion, generating actions by iteratively denoising continuous-valued trajectories [black2024pi_0, intelligence2025pi05, lin2025onetwovla] (see Fig. 1(b)). While expressive, continuous diffusion introduces two fundamental issues. First, it operates in a continuous Euclidean space that is semantically misaligned with the discrete and symbolic structure of pretrained VLM/VLA backbones, weakening the coupling between language instructions and action generation. Second, continuous trajectories are physically inconsistent with real-world robot execution. Hardware constraints such as control frequency, encoder resolution, and actuation latency inherently quantize any commanded continuous signal into coarse, noisy motions, meaning that continuous denoisers learn a mapping that does not reflect how actions are actually executed on real robots. These limitations have inhibited the generalization and fine-grained control capabilities of existing VLA models.
@@
These observations motivate a new perspective: to design an action modeling approach that (i) remains compatible with the symbolic structure of pretrained vision–language backbones, (ii) aligns with the quantized nature of real robot control, and (iii) supports high-resolution action vocabularies that capture fine-grained manipulations. To this end, we propose $\mathcal{E}_{0}$, a continuized discrete diffusion framework that formulates action generation as iterative denoising over Gaussian-noised one-hot action vectors (see Fig. 1(c)). By applying Gaussian noise directly to discrete action embeddings, $\mathcal{E}_{0}$ follows Tweedie’s formulation and maintains forward–reverse consistency, avoiding the distributional mismatch inherent in mask-based diffusion. Unlike AR-based methods, our approach supports arbitrarily fine discretization bins, enabling high-resolution action modeling beyond the constraints of language tokenizers. At the same time, by operating in a discrete space, $\mathcal{E}_{0}$ remains aligned with the symbolic representations used in pretrained VLM/VLA backbones 111see Analysis in Section 6 of the appendix., strengthening semantic conditioning. Furthermore, because discrete diffusion models the true quantized nature of robot control, its Bayes-optimal denoiser captures the correct discrete action distribution and yields stronger generalization than continuous models 222see Analysis in Section 7 of the appendix..
@@
To further enhance robustness, we introduce a spherical viewpoint perturbation augmentation that perturbs input observations based on the relative camera–scene geometry. This augmentation improves cross-view consistency and reduces sensitivity to camera shifts without requiring additional data collection, addressing a long-standing practical challenge.
We evaluate our approach across three distinct simulation environments (LIBERO [liu2023libero], VLABench [zhang2024vlabench], ManiSkill [tao2024maniskill3]) encompassing diverse scenes, object types, and task difficulties, and further validate it on a real-world robotic platform. Experimental results demonstrate that the discrete diffusion-based action generation effectively preserves the semantic understanding of pretrained VLMs while enabling fine-grained action synthesis. $\mathcal{E}_{0}$ achieves state-of-the-art performance across 14 diverse and semantically challenging environments, surpassing baseline models by an average margin of 10.7%. On VLABench [zhang2024vlabench], our model accurately interprets task instructions and subtle linguistic cues, successfully distinguishing between different styles of paintings and grasping poker cards or mahjong tiles with specific patterns. Moreover, it excels at two of the most challenging manipulation tasks—plug insertion and rectangular peg insertion—where most existing models struggle. In addition, our proposed spherical-view perturbation augmentation provides a plug-and-play solution for improving viewpoint generalization without additional data collection or manual effort, yielding consistent performance gains on both baseline models and our proposed $\mathcal{E}_{0}$.
@@
    We propose $\mathcal{E}_{0}$, a continuized discrete diffusion framework for action modeling that supports arbitrarily fine discretization. This formulation enables high-precision action representation while maintaining efficient inference, and preserves compatibility with pretrained vision–language models for accurate perception, grounding, and manipulation.
@@ 2 Related Work
Autoregressive VLA Models. Early efforts such as RT-1 [brohan2022rt] and RT-2 [zitkovich2023rt] pioneered Vision–Language–Action (VLA) models, demonstrating the potential of scaling language-conditioned policies. Building on this foundation, OpenVLA [kim2024openvla] released the first open-source autoregressive VLA by integrating Llama-2 [touvron2023llama] with strong vision encoders (DINOv2 [oquab2023dinov2], SigLIP [zhai2023sigmoid]), while SpatialVLA [qu2025spatialvla] introduced Ego3D position encoding to inject explicit 3D spatial cues. To enhance tokenized action generation, OpenVLA-OFT [kim2025fine] proposed an optimized fine-tuning recipe with parallel decoding and continuous action representations. $\pi_{0}$ FAST [pertsch2025fast] designed a frequency-aware tokenization scheme for efficient autoregressive training. In parallel, reasoning-augmented approaches such as CoT-VLA [zhao2025cot] and generative rollouts (GR-1 [wu2023unleashing], GR-2 [cheang2024gr]) extend VLAs with textual or visual chain-of-thought to guide planning. Our work builds on these advances by preserving the strong contextual alignment afforded by autoregressive decoding, while addressing its inherent limitations in producing fine-grained continuous actions.
@@
Diffusion-based VLA Models. The application of diffusion to robotic imitation learning was first introduced by Diffusion Policy [chi2023diffusion], demonstrating its promise for continuous action generation. Subsequent works have advanced this paradigm through increasingly sophisticated designs: CogACT [li2024cogact] developed diffusion action transformers with favorable scaling properties, while RDT [liu2024rdt] exploited DiT’s cross-modal fusion to generate complete action sequences directly from noisy action chunks, later extended by A0 [xu2025a0] with affordance-based reasoning. Generalist frameworks such as $\pi_{0}$ [black2024pi_0] and $\pi_{0.5}$ [intelligence2025pi05] proposed powerful flow models for broad robotic control, achieving strong performance across diverse real-world tasks. Other recent directions include OneTwoVLA [lin2025onetwovla], which enables dynamic switching between reasoning and execution, GR00t N1 [bjorck2025gr00t] with a dual-system design, and GO-1 [bu2025agibot], a scalable generalist policy using latent actions. Finally, Hybrid VLA [liu2025hybridvla] combines autoregressive and diffusion decoding to leverage their complementary strengths. In contrast, our work preserves the expressive continuous modeling ability of diffusion while maintaining the emergent language grounding and multimodal reasoning capacities of AR VLAs.
@@
Discrete Diffusion. Recent advances have sparked growing interest in discrete diffusion models. A common formulation treats discrete diffusion as a masked modeling problem, in which noise is simulated through random token masking and the model is trained to reconstruct the original input [nie2025large, wu2025fast, ye2025dream]. Recently, an attempt has been made to extend discrete diffusion to VLA models [liang2025discrete]; however, their design follows a BERT-like [devlin2019bert] masking mechanism and requires additional architectural complexity to achieve competitive performance. In contrast, our discrete diffusion provides a reformulation that operates directly on float-encoded one-hot class representations, avoiding both continuous latent diffusion and heuristic masking strategies. During inference, the model predicts a clean categorical state at each step by applying an $\arg\max$ operation over output probabilities and converting the result back into a one-hot vector. Such an autoregressive-style feedback mechanism enables gradual refinement of predictions. Building upon this foundation, our work makes more stable action generation and achieves competitive performance in complex control scenarios.
@@ 3 Method
Figure 3: Benchmarks for evaluation. (a) LIBERO [liu2023libero]: tasks with varying objects, layouts, and goals, including long-horizon settings. (b) ManiSkill [tao2024maniskill3]: diverse fine-grained manipulation skills (push, pick, stack, insert, plug). (c) VLABench [zhang2024vlabench]: open-ended tasks requiring language grounding and commonsense reasoning (select toy/fruit/painting/poker/mahjong).
@@ 3.1 Preliminaries
Forward in DDPM.
As described in [ho2020denoising, kingma2021variational, chan2024tutorial], the objective of the forward process is to produce a sequence of intermediate variables $x_{1},\ldots,x_{T-1}$ from a clean input $x_{0}$ (which may represent actions, images, or other modalities). At each timestep, the perturbation is defined by:
$x_{t}=\mathcal{N}\!\left(x_{t}\,\middle|\,\sqrt{\bar{\alpha}_{t}}\,x_{0},\;(1-\bar{\alpha}_{t})\mathbf{I}\right)\quad t=1,\ldots,T-1,$ (1)
where $\bar{\alpha}_{t}$ denotes the scaling and noise schedules. This forward process requires no training: given $x_{0}$, one can directly run diffusion to obtain the perturbed sequence $x_{1},\ldots,x_{T}$.
Inference in DDPM. Once the denoising model $\hat{\mathbf{x}}_{\theta}$ has been trained, it can be applied for inference. The goal of inference is to generate samples by iteratively drawing from the conditional distributions $p_{\theta}(x_{t-1}\mid x_{t})$ over the sequence of states $x_{T},x_{T-1},\ldots,x_{1}$. Since this corresponds to the reverse diffusion process, the sampling must be performed recursively according to
$x_{t-1}\sim p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\!\left(x_{t-1}\,\middle|\,\mu_{\theta}(x_{t}),\;\sigma^{2}(t)\,\mathbf{I}\right).$ (2)
Here, $\mu_{\theta}(x_{t})$ denotes the predicted denoised mean estimated by the neural network, indicating the model’s estimate of the clean sample at step $t$, while $\sigma^{2}(t)$ represents the time-dependent variance that controls the stochasticity of the reverse diffusion process.
@@ 3.2 Architecture and Pipeline
VLM Backbone and Action Expert.
As shown in Fig. 2(a), we build on PaliGemma [beyer2024paligemma], an open-source VLM, and additionally incorporate a 300M-parameter action expert as the representation backbone. Unlike traditional diffusion-based approaches [black2024pi_0, intelligence2025pi05] that map the outputs of a VLM into continuous action representations via an MLP, our method aligns with autoregressive VLA paradigms by discretizing each action into an arbitrary number of bins. Notably, autoregressive VLAs  [brohan2022rt, zitkovich2023rt, kim2024openvla] typically fix the bin size at 256, where the resolution of action representation is inherently limited by the efficiency of autoregressive decoding. Unlike prior discretization schemes constrained by fixed bin sizes, our method permits an arbitrary number of bins, achieving finer action granularity without sacrificing inference speed.
Consistent with prior work [brohan2022rt, zitkovich2023rt, kim2024openvla, liang2025discrete], we adopt a quantile-based discretization scheme, using the 1st–$N$th percentiles (with $N$ up to 2048 or higher) for each dimension to mitigate the influence of outliers. This effectively filters out extreme values and ensures smooth and stable performance during robotic inference. A single-timestep action can be represented as $D_{a}=7$ (3 for translation, 3 for rotation, and 1 for the gripper). Alternatively, it can also be defined as $D_{a}=8$ ( 7 for joint angles of the manipulator and 1 for the gripper, e.g., Franka Research 3). In our experiments, we set the number of bins to $2048$, which provides a favorable trade-off between representational precision and inference efficiency. For action chunking, tokens from $H$ future timesteps are arranged into a fixed layout, yielding a total of $L_{a}=H\times D_{a}$ action positions. Benefiting from the strong generative capacity of our authentic, discrete diffusion-based action modeling, we are able to set $H=50$ (consistent with [black2024pi_0]), while still maintaining high-quality and coherent action generation.
@@
Visual features $I_{t}^{i}$ and proprioceptive states $q_{t}$ are processed by their respective encoders and linearly projected into the same embedding space as the language tokens, yielding a unified cross-modal representation. Each action $a_{t}$ is discretized using the 1st–$N$th percentiles (with $N$ up to 2048 or higher) to reduce outlier influence, producing a sequence of discrete tokens $\tilde{A}_{t}$ of length $L$, represented as one-hot vectors.
During training, a timestep $\tau\in[0,1]$ is uniformly sampled, and Gaussian noise $\varepsilon\!\sim\!\mathcal{N}(0,I)$ is added to the discretized actions:
$\tilde{A}_{t}^{\tau}=\tau\tilde{A}_{t}+(1-\tau)\varepsilon.$ (3)
Since discretized one-hot actions are inherently sharp, we apply an additional smoothing factor $\alpha=0.1$ on $\tilde{A}_{t}$ before adding noise, which improves training stability.
Following [black2024pi_0], we draw $\tau$ from a beta distribution instead of the uniform schedule [lipman2022flow, liu2022rectified]. This bias toward smaller $\tau$ (i.e., higher-noise regions) reflects the stronger multimodal constraints in embodied action prediction compared to image generation, encouraging the model to learn robust denoising behavior under high uncertainty.
@@
Inference Pipeline. During inference, we begin from an initial noisy action sequence and perform a multi-step iterative denoising procedure. At the first iteration, the observation $o_{t}$ is encoded and processed by the model to produce the initial key–value cache $\mathit{KV}(o_{t})$, which stores the key–value pairs required for cross-attention. This cache is then reused across all subsequent denoising steps, so that only the action tokens need to be recomputed.
At each iteration $i\in\{1,\ldots,N\}$, the model receives the current noisy action sequence $\tilde{A}_{t}^{\tau_{i}}$ along with the fixed observation cache $\mathit{KV}(o_{t})$, and predicts a categorical distribution over discrete action tokens. These tokens are decoded into one-hot representations, which serve as intermediate actions for the next iteration. The forward noise process is then reapplied to obtain the next noisy sequence:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\hat{A}_{t}^{(i)}+(1-\tau_{i+1})\,\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,I).$ (6)
After $N$ iterations, the final discrete tokens are deterministically detokenized into continuous actions to form the reconstructed action chunk $A_{t}=[a_{t,1},\ldots,a_{t,H}]$.

### v2

# via Tweedie Discrete Diffusion

@@ Abstract
Vision–Language–Action (VLA) models offer a unified framework for robotic manipulation by integrating visual perception, language understanding, and control generation. However, existing VLA systems still struggle to generalize across diverse tasks, scenes, and camera viewpoints, and often produce coarse or unstable actions. We argue that these limitations are closely tied to the structural properties of actions in VLA settings, including the inherent multi-peaked nature of action distributions, the token-based symbolic reasoning of pretrained VLM/VLA backbones, and the effective finite resolution imposed by real-world robotic control. Motivated by these properties, we introduce $\mathcal{E}_{0}$, a tweedie discrete diffusion framework that formulates action generation as iterative denoising over quantized action tokens. By operating in a discrete action space with a principled diffusion process, $\mathcal{E}_{0}$ naturally aligns with token-based reasoning, supports fine-grained yet executable action control, and avoids the distributional mismatch of masking-based discrete diffusion. We further introduce a spherical viewpoint perturbation augmentation to enhance robustness to camera shifts without additional data. Experiments on LIBERO, VLABench, ManiSkill, and a real-world Franka arm demonstrate that $\mathcal{E}_{0}$ achieves state-of-the-art performance across 14 diverse environments, outperforming strong baselines by 10.7% on average.
Vision-Language-Action Models, Embodied AI
@@ 1 Introduction
Figure 1: Overview of action modeling paradigms. (a) Discrete modeling: Traditional autoregressive (AR) approaches (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024) and recent mask-based discrete diffusion methods (Liang et al., 2025), which operate over a small discrete action vocabulary. (b) Continuous modeling: Continuous diffusion–based policies (Liu et al., 2024; Xu et al., 2025) and AR–diffusion hybrids (Black et al., 2024; Intelligence et al., 2025; Lin et al., 2025) that regress continuous actions. (c) Our approach: $\mathcal{E}_{0}$ integrates AR-style conditioning with a tweedie discrete diffusion framework, enabling efficient action generation while preserving compatibility with pretrained vision–language backbones and supporting fine-grained action control.
@@
Recent VLA systems have explored a diverse range of formulations for action generation. Representative approaches include autoregressive token prediction over discretized action spaces (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024; Pertsch et al., 2025), masked diffusion over discrete action tokens (Nie et al., 2025; Wu et al., 2025; Ye et al., ), as well as continuous diffusion models that iteratively denoise real-valued action trajectories (Black et al., 2024; Intelligence et al., 2025; Lin et al., 2025). While differing in architectural design and training objectives, these methods largely adopt action representations inherited from existing sequence modeling or diffusion paradigms, rather than being derived from a principled analysis of action features specific to VLA settings.
To better understand the challenges of action modeling in VLA systems, it is useful to examine the structural properties of actions in this setting. First, robotic actions are inherently multi-peaked (Florence and Lynch, 2021; Shafiullah et al., 2022; Florence et al., 2022): under the same task specification and visual context, multiple distinct action sequences may successfully accomplish the task, making action prediction fundamentally a distribution modeling problem rather than single-valued regression (see Appendix K.1 for additional analysis). Second, action generation in VLA systems is tightly coupled with pretrained vision–language backbones, where high-level reasoning, instruction following, and semantic conditioning are mediated through discrete symbolic tokens (Liang et al., 2025; Li et al., 2025; Luo et al., 2026). These backbones are optimized with token-level classification objectives, under which semantic information is primarily encoded through relative feature directions, shaping the representational and optimization characteristics of VLM/VLA models (see Appendix K.2 for additional analysis). Third, although physical execution evolves in continuous time, practical robotic systems operate under finite control rates, encoder resolution, actuation latency, and environmental noise. These factors collectively impose an effective finite resolution on action execution at the policy level, which may differ substantially from nominal hardware specifications due to real-world physical constraints (see Appendix K.4 for additional analysis). Taken together, these properties highlight fundamental constraints on action modeling in VLA systems, motivating formulations that can capture multi-peaked behaviors, align with token-based reasoning, and operate under finite-resolution execution limits.
For multi-peaked action distributions, discrete representations naturally preserve multiple valid modes without collapsing them into a single averaged prediction. For action generation mediated by discrete symbolic tokens in pretrained vision–language backbones, discrete action spaces enable representational and optimization-level alignment with token-based reasoning. Finally, the effective finite resolution imposed by real-world robotic control makes discrete abstractions well matched to the granularity at which actions can be reliably generated and executed. Collectively, these observations suggest that discreteness in action modeling follows naturally from the multi-peaked structure of actions, the token-based reasoning of VLA models, and the finite resolution of real-world control.
@@
To this end, we propose $\mathcal{E}_{0}$, a tweedie discrete diffusion framework that formulates action generation as iterative denoising over Gaussian-noised one-hot action vectors (see Fig. 1(c)). By applying Gaussian noise directly to discrete action spaces, $\mathcal{E}_{0}$ follows Tweedie’s formulation and maintains forward–reverse consistency, avoiding the distributional mismatch inherent in mask-based diffusion (see Appendix L for additional analysis). Unlike AR-based methods, our approach supports arbitrarily fine discretization bins, enabling high-resolution action modeling beyond the constraints of language tokenizers. At the same time, by operating in a discrete space, $\mathcal{E}_{0}$ remains aligned with the symbolic representations used in pretrained VLM/VLA backbones, strengthening semantic conditioning. Furthermore, because discrete diffusion models the true quantized nature of robot control, its Bayes-optimal denoiser captures the correct discrete action distribution and yields stronger generalization than continuous models.
@@
To further enhance robustness, we introduce a spherical viewpoint perturbation augmentation that perturbs input observations based on the relative camera–scene geometry. This improves cross-view consistency and reduces sensitivity to camera shifts without requiring additional data collection, addressing a long-standing practical challenge.
@@
    We propose $\mathcal{E}_{0}$, a tweedie discrete diffusion framework for action modeling that supports arbitrarily fine discretization. This formulation enables high-precision action representation while maintaining efficient inference, and preserves compatibility with pretrained vision–language models for accurate perception, grounding, and manipulation.
@@ 2 Related Work
Autoregressive VLA Models. Early efforts such as RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) pioneered VLA models, demonstrating the potential of scaling language-conditioned policies. Building on this foundation, OpenVLA (Kim et al., 2024) released the first open-source autoregressive VLA by integrating Llama-2 (Touvron et al., 2023) with strong vision encoders (DINOv2 (Oquab et al., 2023), SigLIP (Zhai et al., 2023)), while SpatialVLA (Qu et al., 2025) introduced Ego3D position encoding to inject explicit 3D spatial cues. To enhance tokenized action generation, OpenVLA-OFT (Kim et al., 2025) proposed an optimized fine-tuning recipe with parallel decoding and continuous action representations. $\pi_{0}$ FAST (Pertsch et al., 2025) designed a frequency-aware tokenization scheme for efficient autoregressive training. In parallel, reasoning-augmented approaches such as CoT-VLA (Zhao et al., 2025) and generative rollouts (GR-1 (Wu et al., 2023), GR-2 (Cheang et al., 2024)) extend VLAs with textual or visual chain-of-thought to guide planning. Our work builds on these advances by preserving the strong contextual alignment afforded by autoregressive decoding, while addressing its inherent limitations in producing fine-grained continuous actions.
@@
Diffusion-based VLA Models. The application of diffusion to robotic imitation learning was first introduced by Diffusion Policy (Chi et al., 2023), demonstrating its promise for continuous action generation. Subsequent works have advanced this paradigm through increasingly sophisticated designs: CogACT (Li et al., 2024) developed diffusion action transformers with favorable scaling properties, while RDT (Liu et al., 2024) exploited DiT’s cross-modal fusion to generate complete action sequences directly from noisy action chunks, later extended by A0 (Xu et al., 2025) with affordance-based reasoning. Generalist frameworks such as $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) proposed powerful flow models for broad robotic control, achieving strong performance across diverse real-world tasks. Other recent directions include OneTwoVLA (Lin et al., 2025), which enables dynamic switching between reasoning and execution, GR00t N1 (Bjorck et al., 2025) with a dual-system design, and GO-1 (Bu et al., 2025), a scalable generalist policy using latent actions. Finally, Hybrid VLA (Liu et al., 2025) combines autoregressive and diffusion decoding to leverage their complementary strengths. In contrast, our work preserves the expressive continuous modeling ability of diffusion while maintaining the emergent language grounding and multimodal reasoning capacities of AR VLAs.
@@
Discrete Diffusion. Recent advances have sparked growing interest in discrete diffusion models. A common formulation treats discrete diffusion as a masked modeling problem, in which noise is simulated through random token masking and the model is trained to reconstruct the original input (Nie et al., 2025; Wu et al., 2025; Ye et al., ). Recently, an attempt has been made to extend discrete diffusion to VLA models (Liang et al., 2025); however, their design follows a BERT-like (Devlin et al., 2019) masking mechanism and requires additional architectural complexity to achieve competitive performance. In contrast, our discrete diffusion provides a reformulation that operates directly on float-encoded one-hot class representations, avoiding both continuous latent diffusion and heuristic masking strategies. During inference, the model predicts a clean categorical state at each step by applying an $\arg\max$ operation over output probabilities and converting the result back into a one-hot vector. Such an autoregressive-style feedback mechanism enables gradual refinement of predictions. Building upon this foundation, our work makes more stable action generation and achieves competitive performance in complex control scenarios.
@@ 3.1 Preliminaries
Continuous-Time Flow Matching. Flow matching (Lipman et al., 2022; Tong et al., 2023) formulates generative modeling as learning a continuous-time velocity field that transports samples from a simple noise distribution to a target data distribution. Given a clean sample $x_{0}$ and noise $\epsilon\sim\mathcal{N}(0,I)$, an intermediate point is constructed as
$x_{t}=t\cdot\epsilon+(1-t)\cdot x_{0},\quad t\in[0,1].$ (1)
The model is trained to predict the velocity
$u_{t}=\epsilon-x_{0},$ (2)
using a regression objective.
Inference via ODE Integration. At inference time, samples are generated by solving the ordinary differential equation
$\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t),$ (3)
from $t=1$ to $t=0$ using a numerical solver. This process is deterministic and does not require an explicit noise variance schedule.
@@ 3.2 Architecture and Pipeline
VLM Backbone and Action Expert. As shown in Fig. 2(a), we build on PaliGemma (Beyer et al., 2024), an open-source VLM, and additionally incorporate a 300M-parameter action expert as the representation backbone. Unlike traditional diffusion-based approaches (Black et al., 2024; Intelligence et al., 2025) that map the outputs of a VLM into continuous action representations via an MLP, our method discretizes actions into bins with arbitrary granularity, while recording the corresponding statistics to enable accurate recovery of the original continuous actions. Notably, autoregressive VLAs (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024) typically encode discrete actions by occupying tokens in the original vocabulary and fix the size of 256. Therefore, increasing the action resolution requires expanding the vocabulary, which incurs additional inference overhead and often necessitates retraining the model. In contrast, our method does not rely on the underlying token vocabulary: action discretization is implemented as a numerical mapping, enabling arbitrarily fine action granularity without increasing inference cost or modifying the model architecture.
Consistent with prior work (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024; Liang et al., 2025), we adopt a quantile-based discretization scheme (using the 1st–$M$th percentiles for each dimension to mitigate the influence of outliers). This effectively filters out extreme values and ensures smooth and stable performance during robotic inference. A single-timestep action can be represented as $D_{a}=7$ (3 for translation, 3 for rotation, and 1 for the gripper). Alternatively, it can also be defined as $D_{a}=8$ ( 7 for joint angles and 1 for the gripper, e.g., Franka Research 3). In our experiments, we set the number of bins to $2048$. For action chunking, tokens from $H$ (with $H=50$ in our experiments) future timesteps are arranged into a fixed layout, yielding a total of $L_{a}=H\times D_{a}$ action positions.
Figure 3: Benchmarks for evaluation. (a) LIBERO (Liu et al., 2023): tasks with varying objects, layouts, and goals, including long-horizon settings. (b) ManiSkill (Tao et al., 2024): diverse fine-grained manipulation skills (push, pick, stack, insert, plug). (c) VLABench (Zhang et al., 2024): open-ended tasks requiring language grounding and commonsense reasoning (select toy/fruit/painting/poker/mahjong).
@@
Visual features $I_{t}^{i}$ and proprioceptive states $q_{t}$ are processed by their respective encoders and linearly projected into the same embedding space as the language tokens, yielding a unified cross-modal representation. Each action $a_{t}$ is discretized using the 1st–$M$th percentiles, producing a sequence of discrete tokens $\tilde{A}_{t}$ of length $L$, represented as one-hot vectors.
During training, a timestep $\tau\in[0,1]$ is sampled, and Gaussian noise $\varepsilon\!\sim\!\mathcal{N}(0,I)$ is added to the discretized actions:
$\tilde{A}_{t}^{\tau}=\tau\varepsilon+(1-\tau)\tilde{A}_{t}.$ (4)
Since the discretized one-hot actions are inherently sharp, we apply an additional smoothing factor $\alpha=0.1$ on $\tilde{A}_{t}$ before adding noise, which improves training stability. Following (Black et al., 2024), we draw $\tau$ from a beta distribution instead of the uniform schedule (Lipman et al., 2022; Liu, 2022). This bias toward smaller $\tau$ (i.e., higher-noise regions) reflects the stronger multimodal constraints in embodied action prediction compared to image generation, encouraging the model to learn robust denoising behavior under high uncertainty.
@@
Inference Pipeline. During inference, we initialize the action sequence from Gaussian noise and perform a multi-step iterative refinement process. At the first iteration, the observation $o_{t}$ is encoded once by the model to construct a key–value cache $\mathit{KV}(o_{t})$ for cross-attention. This cache is reused across all subsequent refinement steps, such that only the action tokens are recomputed.
At each iteration $i\in\{1,\ldots,N\}$ (with $N=10$ in our experiments), the model takes as input the current noisy action representation $\tilde{A}_{t}^{\tau_{i}}$ together with the fixed observation cache $\mathit{KV}(o_{t})$, and predicts a categorical distribution over discrete action tokens. The predicted tokens are converted into one-hot representations, which are used to construct an updated action vector for the next iteration. Noise annealing is implicitly controlled by the scalar timestep $\tau_{i}$, which monotonically decreases from $1$ to 0 over refinement steps. Specifically, the next noisy action representation is obtained as:
$\tilde{A}_{t}^{\tau_{i+1}}=\tau_{i+1}\,\varepsilon+(1-\tau_{i+1})\,\hat{A}_{t}^{(i)},\quad\varepsilon\sim\mathcal{N}(0,I),$ (7)
where $\varepsilon$ is fixed-variance isotropic Gaussian noise and $\hat{A}_{t}^{(i)}$ denotes the discrete action estimate at iteration $i$.
After $N$ refinement steps, the final discrete action tokens are deterministically detokenized into continuous actions, yielding the reconstructed action chunk $A_{t}=[a_{t,1},\ldots,a_{t,H}]$. Further details are provided in Appendix A.1.